# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

### v1

@@ 4.1. Overview and Research Questions
The evaluation validates the PRISM framework through three complementary research questions that progress from single-artifact generation with comprehensive constraint enforcement to multi-artifact system generation under scalability stress, culminating in cross-domain transferability assessment where explicit meta-models are unavailable. This progression addresses three fundamental challenges in LLM-driven artifact generation for regulated domains: establishing formal correctness guarantees for individual artifacts, maintaining consistency across interdependent artifacts at system scale, and demonstrating architectural generalizability beyond domains with rich formal specifications.
@@
RQ1 (Single-File Generation)::
Can layered constraint enforcement with evidence-carrying generation achieve both structural correctness and high semantic compliance on single-file AUTOSAR components while converging to valid outputs within one repair iteration?
RQ2 (Multi-File System Generation)::
Can the PRISM framework scale to multi-file AUTOSAR systems while maintaining structural correctness, and what are the dominant failure modes when cross-file dependencies introduce global semantic constraints?
RQ3 (Cross-Domain Transferability)::
Can the UMM-ICM-CVG architecture extend to domains without explicit meta-models through inductive constraint extraction while maintaining layered constraint enforcement and evidence-carrying generation?
@@ 4.2.2. Instance-Constraint Mapping Construction
The resulting ICM transforms natural-language obligations from specification documents into structured, typed constraints aligned with the UMM vocabulary, bridging the gap between informal domain knowledge and formal verification requirements. Applying this process to the AUTOSAR Software Component Template specification (AUTOSAR_TPS_SoftwareComponentTemplate) yielded 1,161 normative constraints governing component architecture, port semantics, interface contracts, and behavioral invariants. The entity linking mechanism successfully anchored 1,045 constraints to verified UMM entities, achieving 90% precision. The remaining constraints flagged for human review primarily involved terminological variations and format inconsistencies in the source specification, which undergo manual validation before inclusion in the final ICM. This precision level demonstrates that the dual-channel extraction approach operates at sufficient reliability for production deployment while maintaining the integrity guarantees required for downstream certification workflows.
@@ 4.3. RQ1: Single-File AUTOSAR Component Generation
RQ1 evaluates the complete PRISM pipeline on single-component AUTOSAR generation, examining whether UMM-ICM-guided retrieval, layered constraint enforcement, and Audit-Guided Repair deliver structural correctness, semantic compliance, and efficient convergence.
@@ 4.3.1. Experimental Setup
RQ1 follows §4.1: single-file AUTOSAR generation with L1+L2 enforcement and one-iteration repair convergence (metrics defined below).
@@
    vLLM (pure prompting baseline): Direct generation without retrieval augmentation or Layer-1 enforcement, mirroring common practice in LLM-based information extraction systems where pretrained models are directly prompted to emit structured records (Dagdelen et al., 2024; Lu et al., 2022).
@@
    vLLM+RAG: Augments prompts with facts retrieved from the AUTOSAR UMM and KG, analogous to instruction-following text-to-database agents that ground extraction in existing schemas (Jiao et al., 2024).
@@
    vLLM+RAG+JSON Schema: Enforces Layer-1 constraints via intermediate JSON instances deterministically projected to ARXML, approximating schema-aligned decoding (Lu et al., 2025).
@@
    vLLM+RAG+GBNF: Applies grammar-constrained decoding using GBNF (GBNF Notation Format) derived from AUTOSAR specifications, masking invalid continuations at each generation step to allow only grammar-admissible prefixes, following the constrained decoding paradigm (Scholak et al., 2021; Beurer-Kellner et al., 2024; Park et al., 2025b).
@@
We measure three dimensions of correctness and efficiency:
@@
[…]
  • Repair efficiency: Average repair iterations required for AGR to converge to valid artifacts.
@@
For semantic validation, we focus on a representative subset of the 1,161 extracted constraints covering mandatory element presence, reference integrity, cardinality restrictions, and type compatibility rules. This subset selection addresses practical constraints: AUTOSAR data used in this study are subject to confidentiality restrictions, and commercial configuration tools limit automated harness construction for the full constraint suite. Integration of the complete constraint set remains ongoing work. Within this tested subset, SHACL and SMT (Satisfiability Modulo Theories) validators cover nearly identical violation patterns; we therefore report SHACL results only.
@@ 4.3.2. Results
Correctness and Repair Convergence.
Table 4 presents structural and semantic correctness rates alongside repair iteration counts. The pure vLLM baseline fails all XSD validation checks, confirming that unconstrained generation produces structurally malformed ARXML. Adding RAG alone achieves 100% XSD compliance through improved context understanding but yields only 30% SHACL compliance with baseline retrieval, indicating that structural correctness does not guarantee semantic constraint satisfaction.
[…]
Repair convergence validates the AGR mechanism: all pipelines with Layer-1 enforcement converge within one iteration on average, compared to two iterations for the unconstrained baseline. For comparison, an LLM-API configuration using the same SHACL validation suite achieves 50% XSD compliance and 10% SHACL compliance, demonstrating the limitations of API-based generation without systematic constraint enforcement.
[…]
Method XSD Pass Rate $_{\text{base}} () SHACL$$_{\text{enh}}$ (%) Average Repair Iterations
vLLM 0 0 2
vLLM+RAG 100 30 1
vLLM+RAG+GBNF 100 30 1
vLLM+RAG+JSON Schema 100 30 100 1
LLM-API (reference) 50 10 1
@@ 4.3.3. Analysis and Discussion
The progression from 0% to 100% XSD compliance (pure prompting to constrained decoding) and from 30% to 100% SHACL compliance (baseline RAG to constraint-aware RAG) validates the core architectural hypothesis of PRISM: structural correctness can be guaranteed at generation time through Layer-1 enforcement, while semantic correctness benefits from explicit constraint materialization during the generation phase rather than solely relying on post-hoc validation.
This finding addresses a fundamental challenge in LLM-driven artifact generation: implicit constraint satisfaction from context alone proves insufficient for domains with rich semantic invariants. The baseline RAG configuration achieves 30% SHACL compliance when the LLM must infer constraints from generic metamodel descriptions, demonstrating inherent limitations of unstructured context. In contrast, constraint-aware retrieval that explicitly tags and injects ICM constraints into prompts enables the LLM to satisfy semantic requirements proactively, reducing AGR to handling residual violations from untested constraint categories.
@@
The 100% audit coverage and structural proof availability for constrained pipelines demonstrate that evidence-carrying generation is practically achievable without prohibitive overhead. The audit trail density of approximately three events per token provides sufficient granularity for root-cause diagnosis during repair while maintaining reasonable storage requirements. The recorded DFA traces enable deterministic reconstruction of generation decisions, supporting both reproducibility requirements in regulated domains and diagnostic capabilities for failed validation.
Repair Efficiency and Convergence Properties.
The consistent one-iteration convergence for pipelines with Layer-1 enforcement empirically validates the bounded repair iterations property discussed in Section 3.6. The key insight underlying this efficiency is dependency-aware prioritization: AGR exploits the constraint hierarchy to repair structural violations before semantic violations, preventing cascading failures that would require multiple repair rounds.
Constraint Subset and Generalization.
The perfect SHACL compliance achieved on the tested constraint subset should be interpreted within its validation scope. The subset covers core constraint categories representative of typical AUTOSAR component generation scenarios—mandatory element presence, reference integrity, cardinality restrictions, and type compatibility. However, extension to the full 1,161-constraint suite requires addressing practical challenges including confidentiality restrictions on AUTOSAR data and tooling limitations for automated validation harness construction. The ongoing integration effort aims to expand coverage while maintaining the demonstrated correctness guarantees.
@@
The latency and token cost results reveal computational trade-offs inherent in constrained generation. Constraint-aware RAG substantially increases input tokens (up to 6,649 tokens in the Full regime) but reduces output tokens through compact, compliant generation. The JSON Schema pipeline achieves lowest latency by exploiting deterministic intermediate representations, while GBNF exhibits higher latency due to per-token automaton state transitions.
Synthesis.
RQ1 validates that the complete PRISM pipeline—encompassing UMM-ICM-guided retrieval, layered constraint enforcement, evidence capture, and AGR—delivers verifiable AUTOSAR component generation with formal correctness guarantees and efficient repair convergence. The results demonstrate that layered constraints, evidence composition, and targeted constraint injection collectively enable trustworthy LLM-assisted software engineering in regulated domains. Critically, this validation encompasses the end-to-end pipeline rather than isolated Layer-1 decoding, confirming that proof-carrying generation with multi-layer verification provides a viable pathway for deploying LLMs in safety-critical artifact generation.
@@ 4.4. RQ2: Multi-File System Generation at Scale
The second research question extends the evaluation from single-component generation (RQ1) to system-level, multi-file AUTOSAR configurations, investigating the feasibility boundaries and architectural challenges that emerge when LLM-driven generation must maintain consistency across multiple interdependent artifacts.
@@ 4.4.1. Experimental Setup
Figure 8. Two-phase AUTOSAR generation, validation, and repair pipeline. Phase 1 synthesizes a high-level blueprint from NLP requirements and AUTOSAR standards, populating the UMM/ICM database. Phase 2 performs constrained LLM decoding to generate component-specific ARXML with JSON Schema enforcement. Post-generation validation applies XSD, SHACL, and SMT checks, with Audit-Guided Repair generating targeted fixes for any violations to produce a verified Decision Package.:
@@
RQ2 tests scalability from single components to multi-file AUTOSAR systems, stressing cross-file references and global invariants that do not arise in RQ1.
@@
We evaluate generated systems across six primary dimensions. File completeness measures whether all blueprint-declared files are produced without truncation, assessed post-generation for all 284 files across the 20 systems as a binary complete-or-incomplete determination. Structural correctness quantifies syntactic compliance with AUTOSAR XSD schemas through xmllint validation, reporting the XSD validation pass rate for all generated ARXML files. Cross-file reference integrity calculates the resolution rate as the proportion of resolvable references over total <REF> elements, measured by parsing all reference targets and verifying their existence in the generated artifact set. Toolchain usability evaluates whether systems successfully import into commercial AUTOSAR configuration tools, specifically Vector DaVinci Developer, yielding a binary pass-or-fail outcome per system. Generation cost captures both token counts and wall-clock time, measured separately for Phase 1 blueprint synthesis and Phase 2 component assembly across all 20 systems.
Expert review follows a structured protocol evaluating four dimensions: requirements alignment assesses whether generated systems satisfy functional specifications and maintain traceability to original requirements; architectural quality evaluates dependency consistency and type-level correctness; engineering quality measures naming conventions, structural modularity, and estimated repair effort; toolchain integration validates whether systems import successfully into commercial AUTOSAR configuration environments. Reviews are conducted by a domain expert with five years of AUTOSAR development experience and aggregated at the group level to identify systematic patterns rather than system-specific idiosyncrasies.
@@ 4.4.2. Results
Structural Correctness and Completeness.
All 20 systems achieve 100% file completeness and 100% XSD validation pass rates across all three complexity tiers, encompassing 284 generated ARXML files (Table 5). Cross-file reference integrity exhibits tier-dependent degradation: Simple systems maintain 100% reference resolution (605 of 605 references resolved), Middle systems decline to 99.4% (1,744 of 1,754 references), and Complex systems reach 92.8% (1,936 of 2,086 references). The overall reference resolution rate across all systems is 96.4% (4,285 of 4,445 references).
@@ 4.4.3. Analysis and Discussion
The structural correctness results presented in Table 5 extend RQ1’s finding that Layer-1 constraint enforcement guarantees structural correctness from single-file to system-level scenarios. Blueprint-guided dual-phase generation with JSON Schema constraint enforcement successfully addresses the syntactic challenges of multi-file artifact generation across all complexity tiers. The blueprint serves three critical functions: establishing globally consistent naming conventions that prevent identifier collisions across files, defining file layout and dependency structure that guides Phase 2 component generation, and providing explicit reference targets that reduce ambiguity during cross-file reference formation.
@@
However, the blueprint’s effectiveness is limited to structural and local semantic properties. While it successfully stabilizes file organization and local type matching (B2 consistently scores 5/5 across all tiers), it fails to enforce global architectural invariants that span multiple components. Expert review findings demonstrate this limitation clearly: dependency consistency (B1) degrades from 5/5 in Simple systems to 2/5 in Middle systems and 1/5 in Complex systems, even as local type matching remains perfect. This divergence indicates that the blueprint captures explicit dependencies declared in component interfaces but fails to encode implicit global constraints such as client-server role separation in diagnostic subsystems or observer-style event propagation in health monitoring architectures.
The Cross-File Reference Boundary.
The progressive degradation pattern observed in Table 5 identifies cross-file reference semantics as the dominant scalability boundary for LLM-driven multi-file generation. This failure mode differs qualitatively from the structural errors observed in RQ1’s unconstrained baseline: rather than producing malformed XML that fails schema validation, the multi-file pipeline generates syntactically valid artifacts that contain semantically inconsistent cross-file references.
Both failure modes—reference hallucination and architectural role drift—stem from limited context windows and independent Phase 2 generation, which preclude global reasoning. Models resolve references using only local context (the current component being generated plus blueprint metadata), lacking the full system view needed to disambiguate similarly-named candidates or detect global pattern violations such as unidirectional data flow or exclusive role assignment.
This finding has important implications for deploying LLM-based generation in multi-file scenarios. While Layer-1 structural constraints prevent local syntactic errors, they cannot alone guarantee global semantic consistency. Addressing the reference resolution boundary requires complementary mechanisms beyond constrained decoding, including explicit reference target canonicalization, stricter interface scoping to reduce choice proliferation, and post-generation validation of system-level invariants before finalization.
Complexity-Dependent Failure Progression.
The tier-specific expert review findings reveal that failure modes intensify with system complexity in predictable patterns. Middle systems encounter isolated role confusion within specific subsystems, affecting localized dependency chains while preserving overall system structure. These failures typically require moderate edits that preserve the generated architectural skeleton. Complex systems exhibit cascading semantic violations that span multiple subsystems, necessitating partial architectural rewrites to restore global invariant satisfaction.
This progression suggests a complexity threshold beyond which purely generative approaches require substantial human refinement. Simple systems with linear topologies and low branching factors remain within the capability envelope of LLM-based generation augmented with blueprint guidance and JSON Schema constraints. Middle systems approach the boundary, achieving high structural quality with localized semantic issues amenable to targeted repair. Complex systems exceed current capabilities for fully autonomous generation, producing structurally sound but semantically inconsistent artifacts that serve as starting points for expert-guided refinement rather than deployment-ready outputs.
Engineering Implications and Mitigation Strategies.
The observed failure modes inform practical deployment strategies for LLM-driven multi-file generation. Three mitigation approaches emerge from the experimental analysis.
First, enhanced blueprint specifications can encode global invariants explicitly. Rather than listing only component interfaces and dependencies, blueprints should specify role assignments determining which components act as clients versus servers in each interaction, communication patterns such as unidirectional data flow and observer relationships, and cross-cutting concerns including security chains and health monitoring topology. This requires extending Phase 1 generation to produce richer architectural specifications that guide Phase 2 generation more tightly.
Second, incremental validation with early feedback can detect semantic violations during Phase 2 generation rather than post-hoc. After generating each component, validating its consistency with previously generated components and the blueprint enables early identification of reference errors and role conflicts. This validation-driven generation approach trades computational cost through multiple validation passes for improved output quality and reduced need for post-generation repair.
Third, hybrid generation with targeted human review acknowledges the complexity boundary identified in this evaluation. For Simple and Middle systems, automated generation with lightweight review focuses on verifying reference resolution and checking for known failure patterns. For Complex systems, generation produces architectural scaffolding that human experts refine to satisfy global invariants, leveraging LLM capabilities for routine structural generation while reserving architectural reasoning for human expertise. This graduated automation model balances efficiency gains from automation with quality requirements for safety-critical deployments.

### v2

@@ 4.1. Overview and Research Questions
The evaluation validates the ATLAS framework through three complementary research questions that progress from single-artifact generation with comprehensive constraint enforcement to multi-artifact system generation under scalability stress, culminating in cross-domain transferability assessment where explicit meta-models are unavailable. This progression addresses three fundamental challenges in LLM-driven artifact generation for regulated domains: establishing formal correctness guarantees for individual artifacts, maintaining consistency across interdependent artifacts at system scale, and demonstrating architectural generalizability beyond domains with rich formal specifications.
@@
RQ1 (Single-File Generation): Can layered constraint enforcement with evidence-carrying generation reliably achieve structural correctness on single-file AUTOSAR components while keeping decoding latency and token cost within practical bounds?
RQ2 (Multi-File System Generation): Can the ATLAS framework scale from single-file to multi-file AUTOSAR systems, preserving Layer-1 structural correctness while assessing how well Layer-2 semantic and Layer-3 logical constraints can be satisfied under realistic cross-file dependencies?
RQ3 (Cross-Domain Transferability): Can the UMM-ICM-CVG architecture, instantiated via the UMM S2 path, transfer to a legal domain without an explicit meta-model and improve legal correctness, citation quality, and rule compliance beyond LLM-only and RAG baselines while maintaining layered constraint enforcement and evidence-carrying generation?
@@ 4.2.2. Instance-Constraint Mapping Construction
The resulting ICM transforms natural-language obligations from specification documents into structured, typed constraints aligned with the UMM vocabulary, bridging the gap between informal domain knowledge and formal verification requirements. This ensures that constraints are not merely text, but executable logic anchored to the domain model.
Constraint Scope and Mapping Statistics.
Applying this dual-channel process to the AUTOSAR Software Component Template specification yielded a total of 1,161 normative constraints. The entity linking mechanism successfully anchored 1,045 of these to verified UMM entities (90% precision), with the remainder flagged for review. For evaluation, we map these constraints into executable L2 validators based on their logical characteristics: 794 SHACL node shapes target graph-structural invariants, and 1,005 SMT assertions target logic and value constraints. Crucially, a significant subset of these constraints governs reference integrity and interaction patterns that, while locally defined, effectively manifest as cross-file dependencies in distributed system architectures. Consequently, satisfying these constraints requires not just local syntactic correctness but global semantic consistency. A subset of constraints unsuitable for formal validation (e.g., purely informational hints) was excluded from the automated pipeline.
@@ 4.3. RQ1: Single-File AUTOSAR Component Generation
RQ1 evaluates the complete ATLAS pipeline on single-component AUTOSAR generation , examining whether UMM-ICM-guided retrieval, layered constraint enforcement, an d audit instrumentation deliver structural correctness with tractable decoding l atency and token overhead.
@@ 4.3.1. Experimental Setup
RQ1 follows §4.1: singl e-file AUTOSAR generation, asking whether Layer-1 constraint enforcement alone c an guarantee structural validity while maintaining acceptable decoding latency a nd token cost.
@@
    vLLM (pure prompting baseline): Direct generation without retrieval augmentation or Layer-1 enforcement, mirroring common practice in LLM-based information extraction systems where pretrained models are directly prompted to emit structured records (Dagdelen2024; Lu2022).
@@
    vLLM+RAG: Augments prompts with facts retrieved from the AUTOSAR UMM and KG, analogous to instruction-following text-to-database agents that ground extraction in existing schemas (Jiao2024).
@@
    vLLM+RAG+JSON Schema: Enforces Layer-1 constraints via intermediate JSON instances deterministically projected to ARXML, approximating schema-aligned decoding (Lu2025b).
@@
    vLLM+RAG+GBNF: Applies grammar-constrained decoding using GBNF (GBNF Notation Format) derived from AUTOSAR specifications, masking invalid continuations at each generation step to allow only grammar-admissible prefixes, following the constrained decoding paradigm (scholakPICARDParsingIncrementally2021; zotero-item-1597; Park2025).
@@
For this structural feasibility study, we focus on three dimensions:
@@
Note: Semantic consistency (SHACL) and logic verification (SMT) metrics are reserved for the system-level evaluation in RQ2, as single-file generation lacks the necessary cross-reference context.
@@ 4.3.2. Results
Structural Correctness (Syntax Feasibility).
Since single-component generation in isolation lacks the system-level context required to satisfy global semantic constraints (e.g., references to external interfaces), RQ1 focuses strictly on structural validity—the prerequisite for any downstream engineering utility. Table 4 presents the XSD validation pass rates. The pure vLLM baseline failed to produce valid XML (0% pass rate), and even the frontier LLM-API baseline (GPT-5) achieved only a 50% pass rate, struggling with the strict nesting and typing rules of ARXML in a zero-shot setting. In contrast, all pipelines equipped with Layer-1 enforcement (JSON Schema or GBNF) achieved 100% structural compliance. This result answers RQ1 affirmatively regarding structural correctness: the proposed L1 masking is sufficient to guarantee that the LLM outputs syntactically valid artifacts, serving as a reliable foundation for the system-level semantic assembly in RQ2.
Table 4. Structural correctness for RQ1: XSD pass rate per pipeline.
Method XSD Pass Rate (%)
vLLM 0
vLLM+RAG 100
vLLM+RAG+GBNF 100
vLLM+RAG+JSON Schema 100
LLM-API 50
@@ 4.3.3. Analysis and Discussion
The stark contrast between the 0% XSD pass rate of the unconstrained vLLM baseline and the 100% pass rate of all constrained pipelines (JSON Schema, GBNF) empirically validates the necessity of Layer-1 enforcement. Notably, even the frontier LLM-API baseline, representing state-of-the-art unconstrained capability, only achieved a 50% pass rate, confirming that probabilistic sampling alone cannot guarantee the strict nesting and typing rules required by ARXML. This finding establishes that Layer-1 structural constraints are not optional optimizations but fundamental prerequisites for reliable artifact generation. By enforcing prefix safety during decoding, ATLAS transforms the LLM from a text generator into a valid model emitter, ensuring that every output is structurally ingestible by downstream engineering tools.
@@
The 100% audit coverage and structural proof availability for constrained pipelines demonstrate that evidence-carrying generation is practically achievable without prohibitive overhead. The audit trail density provides sufficient granularity to trace every structural decision back to the enforcing automaton state. This recorded trace enables deterministic reconstruction of the generation process, satisfying the reproducibility and auditability requirements of regulated domains—a capability entirely absent in standard API-based generation.
@@
The latency and token cost results (Table 2) reveal the inherent trade-offs in constrained generation. Constraint-aware retrieval substantially increases input tokens (up to 6,649 tokens in the Full regime) to provide necessary context, but effectively reduces output tokens through compact, compliant generation. The JSON Schema pipeline achieves the lowest latency by exploiting deterministic intermediate representations, while GBNF exhibits higher latency due to the overhead of per-token automaton state transitions. These metrics suggest that JSON Schema-based enforcement offers the optimal balance between performance and structural guarantees for this domain.
Synthesis: Structural Reliability as a Foundation.
RQ1 validates that ATLAS’s Layer-1 enforcement reliably solves the “syntax problem” of LLM-based generation. While single-file generation inherently precludes the verification of global cross-file semantics (the subject of RQ2), the achievement of 100% structural validity is critical. It ensures that the artifacts passed to the subsequent semantic validation stages are well-formed and parseable. Without this L1 guarantee, any attempt at higher-level semantic reasoning or repair would be thwarted by basic parsing failures, as evidenced by the unconstrained baselines. Thus, RQ1 confirms that the UMM-ICM-CVG architecture provides the necessary structural substrate upon which the complex semantic assembly of RQ2 can be built.
@@ 4.4. RQ2: Multi-File System Generation at Scale
The second research question extends the evaluation from the single-component setting of RQ1 to system-level, multi-file AUTOSAR configurations. Here we stress-test ATLAS under realistic cross-file dependencies, asking whether the UMM-ICM-CVG architecture can maintain the strong Layer-1 guarantees observed in RQ1 while surfacing the semantic and logical failure modes that only emerge at system scale.
@@ 4.4.1. Experimental Setup
Figure 9. Two-phase AUTOSAR generation, validation, and repair pipeline. Phase 1 synthesizes a high-level blueprint from NLP requirements and AUTOSAR standards, populating the UMM/ICM database. Phase 2 performs constrained LLM decoding to generate component-specific ARXML with JSON Schema enforcement. Post-generation validation applies XSD, SHACL, and SMT checks, with Audit-Guided Repair generating targeted fixes for any violations to produce a verified Decision Package.
System architecture for the AUTOSAR experiment shown as a two-phase pipeline. Phase 1 (AUTOSAR Blueprint) consumes NLP requirements, AUTOSAR standards/PNFs, and the AUTOSAR meta-model (XSD/XMI) to synthesize a high-level component and type layout, populating the UMM/ICM database and the AUTOSAR knowledge graph. Phase 2 (Component-wise Generation) assembles a dynamic JSON Schema for each component, performs constrained LLM decoding to emit per-component JSON, and converts that JSON to ARXML. All intermediate artifacts are logged in the Evidence Registry. After generation, only the produced ARXML is validated: it is checked by XSD validation, SHACL semantic validation, and SMT-based logical checks. A Failure Mode Logger and Audit-Guided Repair module analyze any violations to generate targeted fixes, yielding a verified Decision Package for the AUTOSAR Configuration Tool.
@@
RQ2 tests whether ATLAS can scale to multi-file AUTOSAR systems while preserving Layer-1 structural correctness, and to what extent Layer-2 semantic and Layer-3 logical constraints can be satisfied in the presence of thousands of cross-file references and global invariants that do not arise in RQ1.
@@
We evaluate generated systems across six primary dimensions. File completeness measures whether all blueprint-declared files are produced without truncation, assessed post-generation for all 284 files across the 20 systems as a binary complete-or-incomplete determination. Structural correctness quantifies syntactic compliance with AUTOSAR XSD schemas through xmllint validation, reporting the XSD validation pass rate for all generated ARXML files. Cross-file reference integrity calculates the resolution rate as the proportion of resolvable references over total <REF> elements, measured by parsing all reference targets and verifying their existence in the generated artifact set. Toolchain usability evaluates whether systems successfully import into commercial AUTOSAR configuration tools, specifically Etas (ISOLAR-A), yielding a binary pass-or-fail outcome per system. Generation cost captures both token counts and wall-clock time, measured separately for Phase 1 blueprint synthesis and Phase 2 component assembly across all 20 systems.
Expert review follows a structured protocol evaluating four dimensions: requirements alignment assesses whether generated systems satisfy functional specifications and maintain traceability to original requirements; architectural quality evaluates dependency consistency and type-level correctness; engineering quality measures naming conventions, structural modularity, and estimated repair effort; toolchain integration validates whether systems import successfully into commercial AUTOSAR configuration environments.
@@ 4.4.2. Results
Structural Success vs. Semantic Failure.
Table 5 (updated with semantic metrics) illustrates a distinct degradation in model performance as constraints shift from syntactic to semantic and from local to global. Layer-1 (Structure): All 20 systems across all complexity tiers achieved 100% XSD pass rates, confirming that the blueprint-guided constrained decoding effectively solves the syntax generation problem. Layer-2 (Local Semantics via SHACL): When evaluating semantic constraints at the single-file level (checking roughly 794 graph shapes), the pass rates drop precipitously: Simple: 2/57 ($\sim$3.5%), Middle: 16/113 ($\sim$14.2%), and Complex: 6/63 ($\sim$9.5%). These failures largely stem from missing mandatory attributes or incorrect local configurations that are structurally valid but semantically vacuous. Layer-2 (Global Logic via SMT): When imposing SMT constraints that verify cross-file consistency (e.g., end-to-end signal flow, exclusive jurisdiction of components) and stacking them with SHACL, the system-level pass rate was 0% across all tiers. While the model often generated references that resolved to existing IDs (high reference resolution rate, see Table 5), the SMT validation revealed that these connections were frequently semantically invalid (e.g., connecting incompatible ports or creating circular dependencies). This 0% baseline confirms that large-scale system consistency is beyond the zero-shot capability of current LLMs, necessitating the iterative repair and human-in-the-loop workflows ATLAS provides.
@@ 4.4.4. Analysis and Discussion
The universal 100% XSD pass rates across all complexity tiers (Table 5) confirm that blueprint-guided constrained decoding effectively solves the syntax generation problem. The blueprint successfully stabilizes file organization and local type matching (B2 consistently scores 5/5). However, its effectiveness is limited to local properties: expert reviews reveal that dependency consistency degrades significantly in Complex systems (B1: 1/5). This indicates that while blueprints capture explicit interface declarations, they fail to enforce implicit global invariants such as client-server role separation or observer-style propagation.
Fundamentally, the observed 100% structural validity is not merely empirical but theoretical. The Layer-1 constrained decoding effectively functions as a deterministic model transformation that maps a subset of the XSD schema into a strictly enforced finite automaton. Since the decoding process is mathematically bound to valid transitions within this automaton, any structural error would imply a defect in the constraint compilation rules ($\kappa_{\mathrm{struct}}$) rather than a stochastic failure of the LLM. Thus, for syntactic properties, ATLAS transforms the probabilistic generation task into a controllable, correct-by-construction derivation process, leaving only semantic ambiguities to the probabilistic nature of the model.
The Gap Between Reference Existence and Semantic Correctness.
The discrepancy between high reference resolution rates (>>92.8%) and the 0% SMT pass rate highlights a critical hallucination mode: the LLM minimizes “syntax loss” (generating valid-looking IDs) but ignores “logic loss” (semantic coherence). For instance, a SenderPort might syntactically reference a ClientServerInterface, satisfying XML schema rules but violating domain logic. This confirms that reference resolution is a necessary but insufficient proxy for correctness, necessitating the post-hoc L2 verification layer.
Graduated Automation: Design Philosophy.
The observed failure modes validate the “graduated automation” design philosophy. Rather than striving for theoretical 100% zero-shot automation—which proves brittle in complex semantic environments—ATLAS adopts a strategic division of labor. The framework reliably automates the 80% of routine structural construction (validated by 100% XSD compliance), while identifying and surfacing the remaining 20% of complex semantic ambiguities (validated by L2 diagnostics) for targeted repair. The one-iteration convergence of AGR for local issues (SHACL) and the fallback to the Configuration Editor (Route B) for global conflicts (SMT) demonstrate a practical workflow that aligns with industrial safety standards, where human oversight is mandatory but optimized.
Complexity-Dependent Failure Progression and Manual Repair.
The tier-specific expert review findings reveal that failure modes intensify with system complexity. Middle systems encounter isolated role confusion (e.g., client-server violations) that are amenable to automated repair. However, Complex systems exhibit cascading semantic violations—such as broken SecOC authentication chains or inverted health monitoring patterns—that span multiple subsystems. Expert analysis confirms that while automated repair mechanisms (Route A) efficiently resolve local constraint violations, these systemic architectural incoherencies often require human-in-the-loop intervention (Route B). For instance, re-threading a security chain requires high-level architectural intent that current LLMs struggle to infer from local constraints alone. This validates ATLAS’s graduated automation model: the system produces a structurally sound scaffolding, but relies on domain experts to perform the final ”last-mile” semantic repair via the MDE editor, shifting their role from ground-up authoring to high-value architectural correction.
Engineering Implications and Continuous Optimization.
The observed failure modes inform practical deployment strategies. It is important to note that our current experimental setup excluded explicit constraint descriptions from the RAG retrieval context due to the prohibitive volume of constraints linked to individual files. However, ATLAS is designed as a continuous integration system capable of long-term optimization.
First, audit-driven prompt enhancement can be deployed: by analyzing the Evidence Registry to identify frequently violated constraints, the system can selectively inject high-risk constraints into the generation prompt based on file complexity. This dynamic adjustment—analogous to verbal reinforcement learning loops (shinn2023reflexion)—balances constraint adherence against the risks of hallucination and attention drift caused by excessive context, a trade-off that requires long-term iterative tuning (chen2024selfdebug).
Second, learning from golden data offers a pathway to reduce common errors. As the system accumulates verified artifacts (golden data) from successful repairs, fine-tuning the base model can internalize domain invariants that are currently missed. Recent findings suggest that even a small set of high-quality examples can significantly align model behavior with complex domain rules (zhou2023lima).
Nevertheless, our results suggest that even with these enhancements, achieving zero-shot, perfect generation of complex multi-file systems that satisfy all global constraints remains infeasible. The inherent conflict between local generation contexts and global system invariants ensures that the verify-repair loop—and specifically the integration of human expertise for complex ambiguity resolution—will remain an essential architectural component for safety-critical generation.
@@
Graduated Automation: Design Philosophy and Validation.
The observed interaction between structural success and semantic refinement empirically validates the “graduated automation” model. Rather than striving for a theoretical 100% zero-shot automation—which proves brittle in complex semantic environments—ATLAS adopts a strategic division of labor. These results confirm the design goal stated in Section 3: the system successfully automates the routine 80% of configuration tasks (as seen in the structural completeness metrics), allowing engineers to focus their expertise on the remaining 20% of complex semantic disambiguation (as captured in the refinement effort). This split aligns with industrial safety standards, where human oversight is mandatory; ATLAS optimizes this oversight by converting it from manual authoring to high-level evidence review.